OpenStudy (anonymous):

What numbers should be excluded as values of x in the polynomial? x squared minus 4x + 4 over x squared minus 5x + 6

OpenStudy (anonymous):

x cannot be 2 or 3

OpenStudy (aroub):

=x^2-4x+4/x^2-5x+6 =(x-2)(x-2)/(x-3)(x-2) =(x-2)/(x-3) so x is not equal to 3 or 2

OpenStudy (anonymous):

if you notice, the denominator of this beastly fraction can factor into (x-2)(x-3). then if x were equal to 3 or 2, you would be dividing by zero, which is NOT cool....there fore x cannot be 2 or 3...ever.

OpenStudy (anonymous):

oh okay thank you